Beer's Law.
Standard Concentration Absorbance
50 mg/dL 0.150
100 mg/dL 0.300
200 mg/dL 0.600
300 mg/dL 0.900

50. Using the absorbance values and the concentration of each standard complete the appropriate standard curve.
51. If a patient sample was also analyzed at the same time that the standards were analyzed and had an absorbance value of 0.450, what would the glucose concentration of the sample be?
1
Expert's answer
2020-11-20T10:06:53-0500

y = 0.003x

x = y/0.003

c = 0.450/0.003 = 150 mg/dL

the glucose concentration of the sample is 150 mg/dL

Answer: the glucose concentration of the sample is 150 mg/dL